A Bibliography of Publications of George Elmer Forsythe

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

23 February 2024 Version 1.57

Abstract

This bibliography records publications of George E. Forsythe (1917– 1972).

Title word cross-reference

126 [For55c]. 4 [For55c]. Eig3 [For64e]. R(r) < 0 [For41a]. LU [PP73]. QR [PP73]. r [For41a]. s [For68a].

-dimensional [For68a].

13-15 [Cra87]. 15 [FH60b]. 16 [For60a, For60c]. 1959 [Ano59]. 1967 [DH69]. 1968 [Mor69]. 1972 [Dan72]. 1987 [Cra87].

2007 [DFM^{+}07]. 23 [PT53].

50th [DFM^{+}07].

75th [DFM^{+}07].

absolute [For44a]. accelerated [FF52, FF54]. Acceleration [FM51a]. ACM [ACM72, Cra87, For64b, For64c, For64d, For65c, For65d, For65a, For65b, For66b, For66c, For66d, Mis17]. Affix [Cro72]. against [For69c]. aid [FvT62]. Aircraft [FD45, For44a, For46b]. Algebra [DH69, For67a, For68b]. Algebraic [FM69, For53b, FM67, FM71]. algebraicheskikh [FM69]. algebraische [FM71]. ALGOL [FvT62, For64e, BBG^{+}69]. algorithm [FS55b, For60a, For60c, FH60b]. Algorithms [For64a, For66a, Bar02]. Almost [Tau61b]. Alternative [For52a]. altitude [FH56]. Analysis [For59e, PT53, RFW60, Tau63b, For57b, FR58a, FR58b, For58, For59d, For64f, For54b]. Analyst [For53a]. Angeles [HFG51, PT53, For51c]. anniversary [DFM^{+}07]. Annual [ACM72]. Applied [RFW60]. approximations [For56b, For56c]. April [Dan72, Hou73]. arbitrary [For45a]. arrival [DFM^{+}07]. Art [DFM^{+}07]. article [For46a]. articles [Ano72, Ano73]. Ashburn [For46a]. Aspects [DH69]. Astrophysics [Tau63c]. Asymptotic [FM51b, For54a, For55a, For68a]. atmospheric [For47b]. August [ACM72, Mor69, PT53]. Automata [Tau63b]. Automatic [For50b, FW65a, FW65b].

Be [For53b]. Best [FS55a]. Bibliographical [For52b]. Bibliography [For56a, For59a, For4x, For51f, For53c]. binary [For51b]. Biographies [Wei88]. birthday [DFM^{+}07]. Book [Cur78, For60b, For62, For70d, For70e]. books [For56a]. Boston [AFI69, ACM72]. both [For61]. bound [Bar03]. boundary [For50c, For56b, For56c]. bounds [Bar02, For54a, For55a]. Bureau [For51c].

C [Cur78]. California [HFG51, PT53]. Can [For53b]. card [FF52, FF54]. Carlo [HFG51, FL50]. celebrating [DFM^{+}07]. Century [Ewi94]. certain [For54a]. Cesàro [For43]. changing [For47b]. Characteristic [FF51, For50c, FS55b]. chastnykh [VF63]. Chislennoe [FM69]. Classification [For4x, For51f, For53c]. Clenshaw [Bar02]. Collected [Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. Collection [Ano72, Ano73, Boe69]. Comes [For68c]. Communities [Mis17]. comparison [For72a, For72b]. Complex [FH60a]. Computation [Cra87, For70d, For70e, For66a]. Computational [Bau74, DFM^{+}07, For67a, For68b]. Computations [FMM77, FMM80, For56e, Cur78]. Computer [AFI69, Ano59, For57c, FM67, For68c, For69a, For69e, FGH^{+}70, FMM77, Knu96b, FS55b, For56b, For56c, For63, For67b, For70a, Mis17, Nov17, FM69, FM71, FMM80, Knu72, Knu96a, Cur78]. Computer-Verfahren [FM71]. Computers [Tau63b, For56e, For57d, For59b]. computes [For55c].

Computing [DFM^{+}07, For55b, FH60a, Nas90, For61, Mis17]. Condition [FM52, FM51c]. Conditioned [FS55a]. conditions [For50c]. conducted [DH69]. Conference [ACM72, AFI69, Ano59, Cra87, DFM^{+}07]. congress [Mor69]. Connection [For51e, vN51]. constant [For49]. Constrained [For55b]. Constructive [DH69]. Contemporary [FR58a]. continental [FH56]. continued [For70a]. Continuous [Tau62]. Contributions [Tau54]. Convex [For70c, For55a]. Coriolis [For49]. creation [Nov17]. Crout [For60a, For60c]. curriculum [For64f]. Curves [LF73]. Cyclic [FH60a].

Data [AF58, For57c]. Data-Fitting [For57c]. decomposition [For51g, Forxxb]. dedicated [Ano72, Ano73]. Definite [For70c]. Degree [FG65]. Dekker [For69d]. density [FH56]. density-altitude [FH56]. Department [For57b]. derivations [For52a]. Description [BBG^{+}69]. Design [For69b, For69c, Tau63b]. determinants [For51d]. Determination [For44a, PT53, Tau54]. development [Knu72, Knu96a]. developments [For46b]. diagnosis [FvT62]. Difference [For56b, For56c, FW60, VF63]. differences [For71b]. Differential [FW60, Gea71, RFW60, VF63, FR58b, For71b, Forxxa]. differentsialnykh [VF63]. Digital [For57c, FS55b, For56b, For56c]. Digits [For51e, For51c, For51b, vN51]. dimensional [For68a]. Directions [DFM^{+}07, For68a]. discriminate [For69c]. Discussion [For46a]. distinct [For55c]. distributions [For72a, For72b]. Do [For68c]. Dynamic [HFG45].

E. [For46a]. E1398 [FS60]. Edinburgh [Mor69]. education [For57d, For59a, For69a]. Educational [For63, For57b, For67b]. Eigenvalue [For64e, For56b, For56c]. Eigenvalues [PT53, Tau54]. elements [For71b]. elliptic [For71b, Forxxa]. Elmer [Dan72]. Engineering [For61]. Enough [For70d, For70e]. Equation [For67c, For69e, For69f, For45a, For50c, FM51c, FS55b]. Equations [FM52, For53b, FW60, Gea71, PT53, RFW60, Tau54, VF63, For4x, FM4x, FHR51, For51f, FF52, For53c, FF54, FR58b, For71b, Forxxa, Lan52]. equivalent [For44b]. equivalent-potential [For44b]. Ergodic [Tau61b]. Error [Bau74, Bar02, Bar03]. Errors [For50b, For59c, For50a]. escalator [For52a]. essays [Boe69]. evaluation [Bar02, Bar03]. Exact [For49]. exam [For70b]. examinations [For70a]. exams [For70f]. experimentation [For69c]. Experiments [AF58, FF52, FF54]. Extension [FM52, FM51c]. Eyes [Ewi94].

Federation [Mor69]. Fifteen [For53a]. Fifteen-Foot [For53a]. Finite [For56b, For56c, FW60, For60d, For51g, For71b, Forxxb, VF63]. Finite-Difference [FW60, VF63]. Fitting [AF58, For57c]. flow [For45a, For49]. Foot [For53a]. Foreword [For71a]. formulae [For52a]. Forsythe [Cur78, Dan72, DFM^{+}07, Hou73, Ano72, Ano73, Bar02, Her72, Knu72, Knu96a, Mol72, Nov17, Spj72, Var87, Var90]. Forum [For66a]. Fox [For52a]. Frame [FS55b]. frequencies [For54a]. frequency [For55a]. Functions [For70c, Tau61b]. Fundamental [DH69, For55a]. Future [DFM^{+}07].

G [Cur78]. Games [Tau63c]. Gauss [FM51c, For51a, FM52]. Gene [DFM^{+}07]. generalization [For45a]. Generating [Cro72]. Generation [For51c, For51b, For57c, RH72]. Geometric [PP73]. Geometry [Tau62]. George [DFM^{+}07, Ano72, Ano73, Dan72, Her72, Hou73, Knu72, Knu96a, Mol72, Nov17, Var87, Var90]. Gerling [For51a]. German [FM71]. Get [ATK72]. Golub [DFM^{+}07, Spj72]. good [For69c]. gradient [FM4x, FM51a, FM51b, FF52, FF54, For68a]. Grading [FW65a, FW65b]. Grammars [Cro72]. Graphs [Bau74]. Group [Tau61b].

Hall [Cur78]. Harada [RH72]. height [For44a]. Held [HFG51, PT53]. Help [ATK72]. high [For56e, For57d, For59a, For59b]. high-speed [For56e, For59b]. History [Cra87, Nas90]. horizontal [For45a]. Householder [For54b]. HSNC’87 [Cra87]. Hydrodynamics [Tau63c].

IBM [DH69]. IFIP [Mor69]. II [FH60b, Tau61b]. III [Tau63a]. Ill [ATK72]. Ill-Posed [ATK72]. implications [For63]. Improving [FM52, FM51c]. independent [For43]. influence [Var90]. Information [Mor69, Mor69]. Initial [Gea71]. Institute [PT53]. Integration [For50b]. Interesting [For53b]. International [Mor69]. Inversion [FL50, For51g, Forxxb]. Isn’t [For70d, For70e]. Iterations [PP73, Lan52]. Iterative [FHR51]. IV [Tau62].

Jacobi [FH60a]. January [Dan72, Hou73]. Jersey [Cra87]. John [For51e, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. Joint [AFI69, Ano59]. July [HFG51]. June [DH69, HFG51].

K. [RH72]. know [Nov17].

L [For46a]. L. [For46a]. Laboratory [DH69]. Lagrange [For55b]. Language [BBG^{+}69]. Laplacian [For56b, For56c]. latent [For52a]. learn [For61]. lecture [For51e]. Legendre [For51d]. Let [For69c]. letter [For64b, For64c, For64d, For65c, For65d, For65a, For65b, For65e, For66b, For66c, For66d]. level [For45b]. Linear [FM69, FM52, For53b, FM67, For67a, PT53, Tau54, For4x, FM4x, FM51c, FHR51, For51f, FF52, For53c, FF54, For68b, FM71, Lan52]. lineare [FM71]. lineinykh [FM69]. Logic [Tau61a]. lower [For54a, For55a].

M [Cur78]. MA [AFI69]. Machinery [For50b]. Malcolm [Cur78]. March [Ano59, DFM^{+}07]. Masinnye [FMM80]. matematiceskich [FMM80]. Math [For70d, For70e]. Mathematical [Boe69, Cur78, FMM77, FMM80, For52b]. Mathematics [DFM^{+}07, Ewi94, For57b, For57a, RFW60, For56a, For59a, For61, FGH^{+}70]. Matrices [FS55a]. Matrix [FF51, FL50, FH60a, For51g, Forxxb]. Maximum [FH56, For70c]. May [AFI69, Cra87]. Mechanics [Tau61a]. membership [For64b, For64c, For64d, For65c, For65d, For65a, For65b, For65e, For66b, For66c, For66d]. membranes [For54a, For55a]. Memoriam [Dan72]. Memory [Her72, Ano72, Ano73, Mol72]. Meteorology [Tau63c, HFG45]. Method [FL50, FH60a, FM4x, FM51a, FM51b, For68a, For72a, For72b, HFG51]. Methods [Cur78, For59b, FW60, For67a, FMM77, FMM80, VF63, For41a, FS42, For4x, FHR51, For51f, For51g, FF52, For53c, FF54, For56b, For56c, For56d, For56f, For68b, Forxxa, Forxxb]. metody [FMM80, VF63]. Minima [For55b]. minimized [Lan52]. minimized-iterations [Lan52]. Minimum [For70c]. Moler [Cur78]. monographs [For52b]. Monte [FL50, HFG51]. Monthly [Ewi94]. Multipliers [For55b]. must [For61].

National [For51c]. Near [For59e, For58]. Neumann [For51e, For72a, For72b, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. Nonlinear [LF73]. nonviscous [For49]. Nörlund [For47a]. normal [For72a, For72b]. Note [For44b, For50a, For59c, RH72, Spj72]. Notes [FOP^{+}56, For64e]. Numeric [Cra87]. Numerical [DFM^{+}07, For50b, For53a, For54b, FR58b, For59b, For59e, Forxxa, Gea71, PT53, RFW60, Tau63b, For57b, FR58a, For58, For59d, For64f].

Obtaining [FF51]. Off [For50b, For59c, For50a]. Oh [Wei88]. one [For50c]. only [For50c]. operations [For44a]. Operators [Tau63a, Tau61b]. optimum [FM51a, FM51b, For68a]. order [For41a, For51d, For55c]. Ordinary [Gea71]. organized [Mor69]. Orthogonal [AF58, For57c, Bar02]. Other [Tau62, For72a, For72b].

pages [Cur78]. paper [For69d]. Papers [FOP^{+}56, For57a, Knu96b, Cra87, For64e]. parameter [For49]. Parlett [For64e]. Parsers [Cro72]. Partial [FW60, RFW60, VF63, FR58b, For71b, Forxxa]. particle [For49]. Periodic [Tau61b]. Permutations [RH72]. Ph.D. [For70a]. Pioneers [Wei88]. Pitfalls [For70d, For70e]. Pivoting [For60a, For60c]. plane [For49]. polygonal [For54a]. Polynomial [FG65, For70c, Bar02, Bar03]. Polynomials [AF58, For57c, For51d]. Posed [ATK72]. Positive [For70c]. potential [For44b]. Power [PP73]. Preliminary [For50b]. Prentice [Cur78]. Prentice/Hall [Cur78]. presented [Cra87]. President [For64b, For64c, For64d, For65c, For65d, For65a, For65b, For65e, For66b, For66c, For66d]. pressure [For45b]. Princeton [Cra87]. Principal [FH60a]. Principles [For54b]. Problem [ATK72, FS60, For65f]. Problems [For70f, Gea71, For56b, For56c]. Procedure [For64e]. Proceedings [Ano59, HFG51, Cra87, DH69, Mor69, PT53, ACM72]. processing [Mor69]. program [For57b, For59d, For67b]. Programs [FW65a, FW65b]. proizvodnykh [VF63]. projects [For70f]. propagation [For47b]. properties [FM51b]. Publications [RFW60]. Punched [FF52, FF54]. Punched-card [FF52, FF54].

Quadratic [For67c, For69e, For69f, For70c]. qualifying [For70a]. Quantum [Tau61a]. question [For70b].

Radius [For70c]. Random [For41b, For51e, For43, For47a, For51c, For51b, For72a, For72b, vN51]. Raznostnye [VF63]. reconnaissance [FD45, For46b]. reduction [For45b]. references [For56e, For71b]. regularity [FS42]. Relaxation [For51a, For56d, For56f]. Remark [For60c, FH60b]. Remarks [FS42, For69d, For66e]. Report [For50b]. Reprint [For59c]. Research [DH69]. reshenie [FM69]. resheniya [VF63]. Response [RH72]. Review [For54b, For60b, For60d, For62, Cur78]. revolution [For63]. Riesz [For41a]. Rings [Tau63a]. role [For57d, For59d]. rootfinder [FH60b]. roots [For52a]. Rosary [RH72]. Round [For50b]. Round-Off [For50b]. Rounding [Bar02, Bau74, For59c, Bar03, For50a]. Rounding-Off [For59c, For50a]. Rüschlikon [DH69]. Russian [FM69, For52b, For56a, FMM80, VF63].

sampling [For72a, For72b]. Satisfactory [For67c, For69f]. school [For57d, For59a]. Science [Knu96b, For57d, For67b, For69a, FGH^{+}70, For70a, Mis17, Nov17, Knu72, Knu96a]. sciences [Boe69]. Scientific [Cra87, Nas90, For56e, For66a]. Scientist [For68c]. sea [For45b]. Second [For51d, FG65]. Second-Degree [FG65]. Selected [For56e, Knu96b, For51g, Forxxb]. Selfridge [For60d]. semigroups [For55c, For60d]. series [Bar02]. Sets [Tau61a]. shape [For47b]. Shelf [For53a]. Short [FOP^{+}56, For64e]. Simultaneous [PT53]. Singularity [For58, For59e]. sistem [FM69]. society [Mis17]. Solution [FM69, For50c, FS60, For65f, FM67, Lan52, Tau54, FM71]. Solver [For67c, For69f]. Solving [For53b, For69e, For4x, FM4x, FHR51, For51f, For53c, For71b]. Souriau [FS55b]. Speed [For47b, For56e, For59b]. Sphere [FG65, For70c]. Spline [LF73]. Spring [AFI69]. Standards [For51c]. Stanford [DFM^{+}07, DFM^{+}07, For70b, For70a]. State [DFM^{+}07, FR58a]. States [FH56]. Stationary [FG65]. Students [For57a, For61, Var87, Var90]. Study [LF73]. Suggestions [For57a]. Summability [For41b, For41a, For43, For47a]. Summary [For51e]. summation [FS42]. survey [For52b, For59b]. Surveys [RFW60]. SWAC [AF58, For51b, For55c]. Switzerland [DH69]. Syllabi [For70a]. Symposium [HFG51, DH69, PT53]. Systeme [FM71]. Systems [FM69, FM52, FM67, Tau54, FM51c, For51f, For53c, FM71, Lan52]. tables [For45b]. Talking [For57a]. Technical [FOP^{+}56, For64e]. Techniques [For51e, vN51]. telegrapher [For50c]. temperature [For44b]. Tentative [For51f, For53c]. testing [For51c, For51b]. Tests [For64e]. Theorem [DH69, Spj72]. Theory [For51g, Forxxb, PP73, Tau61a, Tau61b, Tau63b, Tau63c]. thermal [For45a]. Till [For68c]. time [For46b]. Today [For67a, For68b]. Topics [Tau62]. trajectories [For49]. Transformation [FM52, FM51c]. Transformations [FF51].

U.C.L.A. [For57b]. undergraduate [For59d, For64f]. unified [Bar03]. Unit [FG65]. United [FH56]. Universal [For45b]. University [DFM^{+}07, For67b, For70a]. untitled [For60b, For62]. uravnenii [FM69, VF63]. Use [AF58, For57c, For56e]. Used [For51e, vN51].

V [For46a, VF63, RFW60, Tau63b]. Value [Gea71, For56b, For56c]. Values [FH60a, FG65]. Variables [For41b, For43, For47a]. Variational [LF73]. Various [For51e, vN51]. Vectorcardiographic [FvT62]. Vectors [FF51]. Verfahren [FM71]. VI [Tau63c]. Volume [RFW60, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. vorticity [For46a]. vycislenij [FMM80].

W [BBG^{+}69]. War [For46b]. War-time [For46b]. waves [For47b]. weather [FD45, For46b]. Weiss [For46a]. Welcoming [For66e]. Well [ATK72]. Western [Ano59]. wind [For44a, For45a]. work [For69c, Var87]. Works [Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c].

zero [For47a]. Zürich [DH69]. Zürich-Rüschlikon [DH69].

References

ACM:1972:PAA

[ACM72] ACM, editor. Proceedings of the ACM Annual Conference, August 1972, Boston. ACM Press, New York, NY 10036, USA, 1972. LCCN TK 7885 A84p 1972. Two volumes.

Ascher:1958:SEU

[AF58] Marcia Ascher and George E. Forsythe. SWAC experiments on the use of orthogonal polynomials for data fitting. Journal of the ACM, 5 (1):9–21, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

AFIPS:1969:ACPa

[AFI69] 1967 Spring Joint Computer Conference, May 14–16, 1969, Boston, MA, volume 34 of AFIPS conference proceedings. AFIPS Press, Montvale, NJ, USA, 1969. LCCN TK7885.A1 J6 1969.

Anonymous:1959:PWJ

[Ano59] Anonymous, editor. Proceedings of the Western Joint Computer Conference, March 3–5, 1959. Institute for Radio Engineers, New York, NY, USA, 1959.

Anonymous:1972:CAD

[Ano72] Anonymous. Collection of articles dedicated to the memory of George E. Forsythe. Communications of the Association for Computing Machinery, 15(8):713–786, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM Student Competition Award Papers.

Anonymous:1973:CAD

[Ano73] Anonymous. Collection of articles dedicated to the memory of George E. Forsythe. SIAM Journal on Numerical Analysis, 10(2):i–xi + 241– 432, 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Anderson:1972:HHI

[ATK72] Donald G. M. Anderson, Joseph Traub, and W. Kahan. How to help an ill-posed problem get well. In ACM [ACM72], page 663. LCCN TK 7885 A84p 1972. The first annual George Forsythe Memorial Lecture, presented by W. Kahan.

Barrio:2002:REB

[Bar02] Roberto Barrio. Rounding error bounds for the Clenshaw and Forsythe algorithms for the evaluation of orthogonal polynomial series. Journal of Computational and Applied Mathematics, 138(2): 185–204, January 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S037704270100382X.

Barrio:2003:URE

[Bar03] R. Barrio. A unified rounding error bound for polynomial evaluation. Advances in Computational Mathematics, 19(4):385–399, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-

9044 (electronic). URL http://link.springer.com/article/10.
1023/A:1024203520270.

Bauer:1974:CGR

[Bau74] F. L. Bauer. Computational graphs and rounding error. SIAM Journal on Numerical Analysis, 11(1):87–96, March 1974. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//www.jstor.org/stable/2156433. In memoriam George Forsythe.

Bauer:1969:AWI

[BBG^{+}69] Henry Bauer, Sheldon Becker, Susan L. Graham, Edwin Satterthwaite, George E. Forsythe, and Robert W. Floyd. Algol W language description. Report CS 110, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1969.

ii + 103 pp. URL https://bitsavers.org/pdf/stanford/cs_
techReports/CS110_ALGOL_W_Revised_Sep69.pdf.

Boehm:1969:MSC

[Boe69] George A. W. Boehm, editor. The Mathematical sciences: a collection of essays. MIT Press, Cambridge, MA, USA, 1969. x + 271 pp. LCCN QA11 .M39. Edited by the National Research Council’s Committee on Support of Research in the Mathematical Sciences (COSRIMS) with the collaboration of George A. W. Boehm.

Crane:1987:HAC

[Cra87] G. E. Crane, editor. HSNC’87: ACM Conference on the History of Scientific and Numeric Computation, conference proceedings: papers presented at the Conference, Princeton, New Jersey, May 13-15, 1987. ACM Press, New York, NY 10036, USA, October 1987. ISBN 0-89791-229-2. LCCN QA76 .A25 1987.

Crowe:1972:GPA

[Cro72] David Crowe. Generating parsers for affix grammars. Communications of the Association for Computing Machinery, 15(8):728–734, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles in honor of George E. Forsythe (ACM Student Competition Award Papers).

Curran:1978:BRB

[Cur78] M. P. J. Curran. Book review: Computer Methods for Mathematical Computations, by G. Forsythe, M. Malcolm, and C. Moler, 1977; 259 pages. (Prentice/Hall, £12.80). The Computer Journal, 21(4):336, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/4/336.2.full.pdf+html.

Daniel:1972:MGE

[Dan72] James W. Daniel. [In memoriam]: George Elmer Forsythe: January 8, 1917–April 9, 1972. Mathematics of Computation, 26(120):872–828, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-

6842 (electronic). URL http://www.jstor.org/stable/2005865;
https://www.ams.org/journals/mcom/1972-26-120/S0025-5718-
72-99098-9/S0025-5718-72-99098-9.pdf.

Du:2007:SSA

[DFM^{+}07] Ding-Zhu Du, Charles Farbat, Walter Murray, Michael Overton, Haesun Park, Michael Saunders, and James Varah, editors. STAN-FORD 50: State of the Art & Future Directions of Computational Mathematics & Numerical Computing: A conference celebrating the 50th anniversary George Forsythe’s arrival at Stanford and Gene Golub’s 75th birthday, Stanford University, March 29–31, 2007. ????, ????, 2007. ISBN ???? LCCN ????

Dejon:1969:CAF

[DH69] Bruno Dejon and Peter Henrici, editors. Constructive Aspects of the Fundamental Theorem of Algebra: proceedings of a symposium conducted at the IBM Research Laboratory, Zürich-Rüschlikon, Switzerland, June 5–7, 1967. John Wiley and Sons, Inc., New York, NY, USA, 1969. ISBN 0-471-20300-9. LCCN QA212 .C65.

Ewing:1994:CMT

[Ewi94] John Ewing, editor. A Century of Mathematics Through the Eyes of the Monthly. Mathematical Association of America, Washington, DC, USA, 1994. ISBN 0-88385-459-7. xi + 323 pp. LCCN QA7 .C38 1994.

Forsythe:1945:AWR

[FD45] George E. Forsythe and R. B. Doremus. Aircraft weather reconnaissance. Report 105–128-1, Headquarters, Army Air Forces Weather Division, September 1945. 218 pp.

Feller:1951:NMT

[FF51] William Feller and George E. Forsythe. New matrix transformations for obtaining characteristic vectors. Quarterly of Applied Mathematics, 8(4):325–331, 1951. CODEN QAMAAY. ISSN 0033-569x (print), 1552-4485 (electronic).

Forsythe:1952:PCE

[FF52] A. I. Forsythe and G. E. Forsythe. Punched-card experiments with accelerated gradient methods for linear equations. NBS Report 1643, National Bureau of Standards, Los Angeles, CA, USA, March 1952. 29 pp.

Forsythe:1954:PCE

[FF54] A. I. Forsythe and G. E. Forsythe. Punched-card experiments with accelerated gradient methods for linear equations. In Taussky [Tau54], pages 55–69.

Forsythe:1965:SVS

[FG65] George E. Forsythe and Gene H. Golub. On the stationary values of a second-degree polynomial on the unit sphere. Journal of the Society for Industrial and Applied Mathematics, 13(4):1050–1068, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X

(electronic). URL http://www.jstor.org/stable/2946425. Cited

in Åke Björck’s bibliography on least squares, which is available by

anonymous ftp from math.liu.se in pub/references.

Forsythe:1970:CSM

[FGH^{+}70] G. E. Forsythe, B. A. Galler, J. Hartmanis, A. J. Perlis, and J. F. Traub. Computer science and mathematics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(4):19–29, September/October 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Forsythe:1956:MDA

[FH56] George E. Forsythe and Morris S. Hendrickson. Maximum densityaltitude in the continental United States. Bull. Amer. Meteorol. Soc., 27(??):576–579, 1956.

Forsythe:1960:CJM

[FH60a] G. E. Forsythe and P. Henrici. The cyclic Jacobi method for computing the principal values of a complex matrix. Transactions of the American Mathematical Society, 94(??):1–23, 1960. CODEN TAM-TAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Forsythe:1960:RAR

[FH60b] George E. Forsythe and John G. Herriot. Remark on Algorithm 15: rootfinder II. Communications of the Association for Computing Machinery, 3(11):602, November 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1951:IMS

[FHR51] George E. Forsythe, Magnus R. Hestenes, and J. Barkley Rosser. Iterative methods for solving linear equations. Bulletin of the American Mathematical Society, 57(6):480, November 1951. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL

https://www.ams.org/journals/bull/1951-57-06/.

Forsythe:1950:MIM

[FL50] George E. Forsythe and Richard A. Leibler. Matrix inversion by a Monte Carlo method. Mathematical Tables and Other Aids to Computation, 4(31):127–129, July 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL https://www.ams.org/

journals/mcom/1950-04-031/S0025-5718-1950-0038138-X. Cor-

rection published in volume 5, p. 55 (1951).

Forsythe:194x:GMS

[FM4x] George E. Forsythe and Theodore S. Motzkin. On a gradient method of solving linear equations. ???? ????, National Bureau of Standards, Los Angeles, CA, USA, ???? 194x. Undated, and found cited only in [Lan52, ref. 7].

Forsythe:1951:AOG

[FM51a] G. E. Forsythe and T. S. Motzkin. Acceleration of the optimum gradient method. Bulletin of the American Mathematical Society, 57(4):304–305, July 1951. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL https://www.

ams.org/journals/bull/1951-57-04/S0002-9904-1951-09523-3/

S0002-9904-1951-09523-3.pdf. Preliminary report. Abstract 57-4-392.

Forsythe:1951:APO

[FM51b] G. E. Forsythe and T. S. Motzkin. Asymptotic properties of the optimum gradient method. Bulletin of the American Mathematical Society, 57(3):183, May 1951. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL https://www.

ams.org/journals/bull/1951-57-03/S0002-9904-1951-09482-3/
S0002-9904-1951-09482-3.pdf. Abstract 57-3-231.

Forsythe:1951:EGT

[FM51c] George E. Forsythe and Theodore S. Motzkin. An extension of Gauss’s transformation for improving the condition of systems of linear equation. Bulletin of the American Mathematical Society, 57(6):481–581, November 1951. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-

881X (electronic). URL https://www.ams.org/journals/bull/
1951-57-06/S0002-9904-1951-09546-4/S0002-9904-1951-09546-
4.pdf.

Forsythe:1952:EGT

[FM52] George E. Forsythe and Theodore S. Motzkin. An extension of Gauss’ transformation for improving the condition of systems of linear equations. Mathematical Tables and Other Aids to Computation, 6(37):9–17, January 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL https://www.ams.

org/journals/mcom/1952-06-037/S0025-5718-1952-0048162-0/
S0025-5718-1952-0048162-0.pdf.

Forsythe:1967:CSL

[FM67] George E. Forsythe and Cleve B. Moler. Computer Solution of Linear Algebraic Systems. Prentice-Hall, Inc., Englewood Cliffs, NJ 07632, USA, 1967. xi + 148 pp. LCCN QA297 .F57 1967.

Forsait:1969:CRS

[FM69] Dž. Forsaĭt and K. Moler. Chislennoe reshenie sistem lineinykh algebraicheskikh uravnenii. (Russian) [Computer Solution of Linear Algebraic Systems]. Izdatel’stvo Mir, Moscow, USSR, 1969. 167 pp. Translated from the English by V. P. Il′in and Ju. I. Kuznecov. Edited by G. I. Marčuk.

Forsythe:1971:CVL

[FM71] George E. (George Elmer) Forsythe and Cleve B. Moler. Computer-Verfahren für lineare algebraische Systeme. (German) [Computer solution of linear algebraic systems]. Verfahren der Datenverarbeitung. R. Oldenbourg, München, Germany, 1971. ISBN 3-486-33601-0. 168 pp. LCCN ???? Translated from the American original by Christine und Helmuth Späth.

Forsythe:1977:CMM

[FMM77] George E. (George Elmer) Forsythe, Michael A. Malcolm, and Cleve B. Moler. Computer Methods for Mathematical Computations. Prentice-Hall series in automatic computation. Prentice-Hall, Inc., Englewood Cliffs, NJ 07632, USA, 1977. ISBN 0-13-165332-6. xi + 259 pp. LCCN QA297 .F5681. US$16.95. Cited in Åke Björck’s bibliography on least squares, which is available by anonymous ftp

from math.liu.se in pub/references.

Forsythe:1980:MMM

[FMM80] George E. Forsythe, Michael A. Malcolm, and Cleve B. Moler. Mašinnye metody matematičeskich vyčislenij. (Russian) [Computer Methods for Mathematical Computations]. Izdatel’stvo Mir, Moscow, USSR, 1980. 280 pp. Translated from the English by Kh. D. Ikramov.

Forsythe:1956:TNS

[FOP^{+}56] George E. Forsythe, Morio Onoe, Z. Pawlak, R. A. Rubenstein, Marjorie Huse, Stefan Machlup, and Hugh E. Stelson. Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 10(53):25–34, January 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL https://www.ams.

org/journals/mcom/1956-10-053/S0025-5718-1956-0076464-4/
S0025-5718-1956-0076464-4.pdf.

Forsythe:1941:RSM

[For41a] G. E. Forsythe. Riesz summability methods of order r, for R(r) < 0. Duke Mathematical Journal, 8(??):346–349, 1941. CODEN DUM-JAO. ISSN 0012-7094 (print), 1547-7398 (electronic). URL http://

projecteuclid.org/euclid.dmj/1077492650.

Forsythe:1941:SRV

[For41b] George E. Forsythe. Summability of Random Variables. Ph.D. dissertation, Brown University, Providence, RI, USA, 1941. iv +

109 pp. URL https://www.proquest.com/dissertations-theses/
summability-random-variables/docview/301889171.

Forsythe:1943:CSI

[For43] G. E. Forsythe. Cesàro summability of independent random variables. Duke Mathematical Journal, 10(??):397–428, 1943. CODEN DUM-JAO. ISSN 0012-7094 (print), 1547-7398 (electronic). URL http://

projecteuclid.org/euclid.dmj/1077471948.

Forsythe:1944:DAH

[For44a] George E. Forsythe. Determination of absolute height and wind for aircraft operations. Report 708, Headquarters, Army Air Forces Weather Division, June 1944. 69 pp.

Forsythe:1944:NEP

[For44b] George E. Forsythe. Note on equivalent-potential temperature. Bull. Amer. Meteorol. Soc., 25(??):149–151, ???? 1944. See also remarks by Neamtan, pp. 228–229.

Forsythe:1945:GTW

[For45a] George E. Forsythe. A generalization of the thermal wind equation to arbitrary horizontal flow. Bull. Amer. Meteorol. Soc., 26(??):371–375, ???? 1945.

Forsythe:1945:UTR

[For45b] George E. Forsythe. Universal tables for reduction of pressure to sea level. Report 972, Headquarters, Army Air Forces Weather Division, 1945.

Forsythe:1946:DVA

[For46a] George E. Forsythe. Discussion of E. V. Ashburn and L. L. Weiss’s article on vorticity. Trans. Amer. Geophys. Union, 27(??):279–282, ???? 1946.

Forsythe:1946:WTD

[For46b] George E. Forsythe. War-time developments in aircraft weather reconnaissance. Bull. Amer. Meteorol. Soc., 27(??):160–163, ???? 1946.

Forsythe:1947:NSR

[For47a] George E. Forsythe. On Nörlund summability of random variables to zero. Bulletin of the American Mathematical Society, 53(??): 302–313, 1947. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-

881X (electronic). URL http://projecteuclid.org/euclid.bams/
1183510599.

Forsythe:1947:SPA

[For47b] George E. Forsythe. Speed of propagation of atmospheric waves with changing shape. J. Meteorol., 4(??):67–69, ???? 1947.

Forsythe:1949:EPT

[For49] George E. Forsythe. Exact particle trajectories for nonviscous flow in a plane with a constant Coriolis parameter. J. Meteorol., 6(??): 337–346, ???? 1949.

Forsythe:194x:CBM

[For4x] George E. Forsythe. Classification and bibliography of methods of solving linear equations. Report ????, ????, ????, ???? 194x. Undated, and found cited only in [Lan52, ref. 5].

Forsythe:1950:NRE

[For50a] George E. Forsythe. Note on rounding-off errors. U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1950. 3 pp.

Forsythe:1950:REN

[For50b] George E. Forsythe. Round-off errors in numerical integration on automatic machinery — preliminary report. Bulletin of the American Mathematical Society, 56(??):61–??, ???? 1950.

Forsythe:1950:STE

[For50c] George E. Forsythe. Solution of the telegrapher’s equation with boundary conditions on only one characteristic. Journal of Research of the National Bureau of Standards (1934), 44:89–102, 1950. CO-DEN ???? ISSN 0160-1741.

Forsythe:1951:GGR

[For51a] George E. Forsythe. Gauss to Gerling on relaxation. Mathematical Tables and Other Aids to Computation, 5(36):255–258, 1951. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

URL https://www.ams.org/journals/mcom/1951-05-036/S0025-
5718-51-99414-8/S0025-5718-51-99414-8.pdf. Translation, with

notes, of a letter by Gauss.

Forsythe:1951:GTRb

[For51b] George E. Forsythe. Generation and testing of 1,217,370 ‘random’ binary digits on the SWAC. Bulletin of the American Mathematical Society, 57(4):304, July 1951. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL https://www.

ams.org/journals/bull/1951-57-04/S0002-9904-1951-09523-3/
S0002-9904-1951-09523-3.pdf. Abstract only.

Forsythe:1951:GTRa

[For51c] George E. Forsythe. Generation and testing of random digits at the National Bureau of Standards, Los Angeles. Applied mathematics series / National Bureau of Standards, 12(??):34–35, ???? 1951. ISSN 1049-4685.

Forsythe:1951:SOD

[For51d] George E. Forsythe. Second order determinants of Legendre polynomials. Duke Mathematical Journal, 18(??):361–371, 1951. CODEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic). URL

http://projecteuclid.org/euclid.dmj/1077476566.

Forsythe:1951:SJN

[For51e] George E. Forsythe. Summary of John von Neumann’s lecture, various techniques used in connection with random digits. Applied mathematics series / National Bureau of Standards, 12(??):36–38, ???? 1951. ISSN 1049-4685.

Forsythe:1951:TCM

[For51f] George E. Forsythe. Tentative classification of methods and bibliography on solving systems of linear equations. INA Report 52-7, National Bureau of Standards, Los Angeles, CA, USA, 1951. 78 pp.

Forsythe:1951:TSM

[For51g] George E. Forsythe. Theory of selected methods of finite matrix inversion and decomposition. Report 52-5, Institute for Numerical Analysis, National Bureau of Standards, Los Angeles, CA, USA, 1951. 93 pp.

Forsythe:1952:ADF

[For52a] George E. Forsythe. Alternative derivations of Fox’s escalator formulae for latent roots. Quarterly Journal of Mechanics and Applied Mathematics, 5(??):191–195, 1952. CODEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Forsythe:1952:BSR

[For52b] George E. Forsythe. Bibliographical survey of Russian mathematical monographs, 1930–1951. Report 1628, National Bureau of Standards, Los Angeles, CA, USA, March 25, 1952. Supplement in Report 1628a, December 12, 1952, 17 pages.

Forsythe:1953:NAF

[For53a] George E. Forsythe. A numerical analyst’s fifteen-foot shelf. Mathematical Tables and Other Aids to Computation, 7(44):221–228, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853

(electronic). URL https://www.ams.org/journals/mcom/1953-07-
044/S0025-5718-53-99356-9/S0025-5718-53-99356-9.pdf.

Forsythe:1953:SLA

[For53b] George E. Forsythe. Solving linear algebraic equations can be interesting. Bulletin of the American Mathematical Society, 59(4):299– 329, July 1953. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-

881X (electronic). URL http://projecteuclid.org/euclid.bams/
1183518018;   https://www.ams.org/journals/bull/1953-59-
04/S0002-9904-1953-09718-X/.

Forsythe:1953:TCM

[For53c] George E. Forsythe. Tentative classification of methods and bibliography on solving systems of linear equations. In Paige and Taussky [PT53], pages 1–28. LCCN QA3 .U5 no. 29.

Forsythe:1954:ALB

[For54a] George E. Forsythe. Asymptotic lower bounds for the frequencies of certain polygonal membranes. Pacific Journal of Mathematics, 4(3): 467–480, 1954. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-

5844 (electronic). URL http://projecteuclid.org/euclid.pjm/
1103044799.

Forsythe:1954:RHP

[For54b] George E. Forsythe. Review of Householder, Principles of Numerical Analysis. Bulletin of the American Mathematical Society, 60(5):488– 491, ???? 1954. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-

881X (electronic). URL http://projecteuclid.org/euclid.bams/
1183519065.

Forsythe:1955:ALB

[For55a] George E. Forsythe. Asymptotic lower bounds for the fundamental frequency of convex membranes. Pacific Journal of Mathematics, 5 (supplement 1):691–702, 1955. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http://projecteuclid.org/

euclid.pjm/1171984829.

Forsythe:1955:CCM

[For55b] George E. Forsythe. Computing constrained minima with Lagrange multipliers. Journal of the Society for Industrial and Applied Mathematics, 3(4):173–178, December 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Forsythe:1955:SCD

[For55c] George E. Forsythe. SWAC computes 126 distinct semigroups of order 4. Proceedings of the American Mathematical Society, 6:443–447, 1955. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Forsythe:1956:BRM

[For56a] George E. Forsythe. Bibliography of Russian mathematics books. Chelsea Publishing Co., New York, NY, USA, 1956. 106 pp.

Forsythe:1956:DMDa

[For56b] George E. Forsythe. Difference methods on a digital computer for Laplacian boundary value and eigenvalue problems. I. finite difference approximations. Communications on Pure and Applied Mathematics (New York), 9(3):425–434, 1956. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Forsythe:1956:DMDb

[For56c] George E. Forsythe. Difference methods on a digital computer for Laplacian boundary value and eigenvalue problems. I. finite difference approximations. In N. Aronszajn, A. Douglis, and C. B. Morrey, Jr., editors, Transactions of the Symposium on Partial Differential Equations held at the University of California, Berkeley, June 20– July 1, 1955, pages 127–136. Interscience, New York, NY, USA, 1956. Reprint of Communications on Pure and Applied Mathematics 9(3) 1956 with different page numbering.

Forsythe:1956:RM

[For56d] George E. Forsythe. Relaxation methods. In I. S. Sokolnikoff, editor, Mathematical Theory of Elasticity, pages 454–465. McGraw-Hill, New York, NY, USA, second edition, 1956. LCCN ???? Section 125.

Forsythe:1956:SRU

[For56e] George E. Forsythe. Selected references on use of high-speed computers for scientific computations. Mathematical Tables and Other Aids to Computation, 10(53):25–27, January 1956. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

URL https://www.ams.org/journals/mcom/1956-10-053/S0025-
5718-1956-0076464-4/S0025-5718-1956-0076464-4.pdf.

Forsythe:1956:WRM

[For56f] George E. Forsythe. What are relaxation methods? In E. F. Beckenbach, editor, Modern Mathematics for the Engineer, pages 428–447. McGraw-Hill, New York, NY, USA, 1956. LCCN ????

Forsythe:1957:SST

[For57a] G. E. Forsythe. Suggestions to students on talking about mathematics papers. American Mathematical Monthly, 64:16–18, 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). Reprinted in [Ewi94].

Forsythe:1957:EPN

[For57b] George E. Forsythe. The educational program in numerical analysis of the Department of Mathematics, U.C.L.A. In Preston C. Hammer, editor, The Computing Laboratory in the University, pages 145–151. University of Wisconsin Press, Madison, WI, USA, 1957. LCCN ????

Forsythe:1957:GUO

[For57c] George E. Forsythe. Generation and use of orthogonal polynomials for data-fitting with a digital computer. Journal of the Society for Industrial and Applied Mathematics, 5(2):74–88, June 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Forsythe:1957:RCH

[For57d] George E. Forsythe. The role of computers in high school science education. Computers and Automation, 6(??):15–16, August 1957. CODEN CPAUAJ. ISSN 0010-4795, 0887-4549.

Forsythe:1958:SNS

[For58] George E. Forsythe. Singularity and near singularity in numerical analysis. American Mathematical Monthly, 65:229–240, 1958. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Forsythe:1959:BHS

[For59a] George E. Forsythe. Bibliography on high school mathematics education. Computers and Automation, 8(??):17–19, May 1959. CODEN CPAUAJ. ISSN 0010-4795, 0887-4549.

Forsythe:1959:NM

[For59b] George E. Forsythe. Numerical methods for high-speed computers — a survey. In Anonymous [Ano59], pages 249–254.

Forsythe:1959:RNR

[For59c] George E. Forsythe. Reprint of a note on rounding-off errors. SIAM Review, 1(1):66–67, January 1959. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Forsythe:1959:RNA

[For59d] George E. Forsythe. The role of numerical analysis in an undergraduate program. American Mathematical Monthly, 66(8):651–662, October 1959. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Forsythe:1959:SNS

[For59e] George E. Forsythe. Singularity and near singularity in numerical analysis. American Mathematical Monthly, 65(4):229–240, 1959. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Forsythe:1960:ACP

[For60a] George E. Forsythe. Algorithm 16: Crout with pivoting. Communications of the Association for Computing Machinery, 3(9):507–508, September 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1960:BRB

[For60b] George E. Forsythe. Book review: [untitled]. Journal of the American Statistical Association, 55(290):391–392, June 1960. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2281764.

Forsythe:1960:RAC

[For60c] George E. Forsythe. Remark on Algorithm 16: Crout with pivoting. Communications of the Association for Computing Machinery, 3(10): 540, October 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1960:RSF

[For60d] George E. Forsythe. Review of Selfridge, On Finite Semigroups. Mathematics of Computation, 14(70):204–207, April 1960. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/journals/mcom/1960-14-070/S0025-
5718-60-99235-8/S0025-5718-60-99235-8.pdf.

Forsythe:1961:ESM

[For61] George E. Forsythe. Engineering students must learn both computing and mathematics. J. Eng. Educ., 52(??):177–188, 1961.

Forsythe:1962:BRB

[For62] George E. Forsythe. Book review: [untitled]. Journal of the American Statistical Association, 57(300):924–926, December 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281829.

Forsythe:1963:EIC

[For63] George E. Forsythe. Educational implications of the computer revolution. In W. F. Freiberger and William Prager, editors, Applications of Digital Computers, chapter 12, pages 166–178. Ginn, Boston, MA, USA, 1963. LCCN QA76.5.

Forsythe:1964:A

[For64a] G. E. Forsythe. Algorithms. Communications of the Association for Computing Machinery, 7(6):347–349, June 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1964:PLAa

[For64b] George E. Forsythe. President’s letter to the ACM membership. Communications of the Association for Computing Machinery, 7(8):507– 509, August 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1964:PLAb

[For64c] George E. Forsythe. President’s letter to the ACM membership. Communications of the Association for Computing Machinery, 7(10):633– 636, October 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1964:PLAc

[For64d] George E. Forsythe. President’s letter to the ACM membership. Communications of the Association for Computing Machinery, 7(12):697– 699, December 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1964:TPA

[For64e] George E. Forsythe. Tests of Parlett’s ALGOL eigenvalue procedure Eig3 (in Technical Notes and Short Papers). Mathematics of Computation, 18(87):486–487, July 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

https://www.ams.org/journals/mcom/1964-18-087/S0025-5718-
1964-0165669-4/S0025-5718-1964-0165669-4.pdf.

Forsythe:1964:UCN

[For64f] George E. Forsythe. An undergraduate curriculum in numerical analysis. Communications of the Association for Computing Machinery, 7 (4):214–215, April 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1965:PLAc

[For65a] George E. Forsythe. President’s letter to the ACM membership. Communications of the Association for Computing Machinery, 8(10):591, October 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1965:PLAd

[For65b] George E. Forsythe. President’s letter to the ACM membership. Communications of the Association for Computing Machinery, 8(12):727, December 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1965:PLAa

[For65c] George E. Forsythe. President’s letter to the ACM membership: why ACM? Communications of the Association for Computing Machinery, 8(3):143–144, March 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1965:PLAb

[For65d] George E. Forsythe. President’s letter to the ACM membership: why ACM? Communications of the Association for Computing Machinery, 8(7):422–426, July 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1965:PLM

[For65e] George E. Forsythe. President’s letter to the membership. Communications of the Association for Computing Machinery, 8(9):541, September 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1965:SP

[For65f] George E. Forsythe. Solution to problem 5334. American Mathematical Monthly, 72(??):1030, November 1965. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Forsythe:1966:FAA

[For66a] George E. Forsythe. A forum on algorithms: Algorithms for scientific computation. Communications of the Association for Computing Machinery, 9(4):255–256, April 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1966:PLAa

[For66b] George E. Forsythe. President’s letter to the ACM membership. Communications of the Association for Computing Machinery, 9(1):1, January 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1966:PLAb

[For66c] George E. Forsythe. President’s letter to the ACM membership. Communications of the Association for Computing Machinery, 9(4):244, April 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1966:PLAc

[For66d] George E. Forsythe. President’s letter to the ACM membership. Communications of the Association for Computing Machinery, 9(5):325, May 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1966:WR

[For66e] George E. Forsythe. Welcoming remarks. Communications of the Association for Computing Machinery, 9(3):137–138, March 1966. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1967:TCM

[For67a] George E. Forsythe. Today’s computational methods of linear algebra. SIAM Review, 9(3):489–515, July 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). Reprinted in Studies in Numerical Analysis 1, SIAM, Philadelphia, 1968.

Forsythe:1967:UEP

[For67b] George E. Forsythe. A university’s educational program in computer science. Communications of the Association for Computing Machinery, 10(1):3–11, January 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1967:WSQ

[For67c] George E. Forsythe. What is a satisfactory quadratic equation solver? Technical report, Stanford University, Department of Computer Science, Stanford, CA, USA, 1967. 13 pp.

Forsythe:1968:ADD

[For68a] George E. Forsythe. On the asymptotic directions of the s-dimensional optimum gradient method. Numerische Mathematik, 11(1):57–76, January 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Forsythe:1968:TCM

[For68b] George E. Forsythe. Today’s computational methods of linear algebra. In Studies numer. Analysis 1, Sympos. numer. Analysis Iowa 1966, pages 106–132. ????, ????, 1968.

Forsythe:1968:WDT

[For68c] George E. Forsythe. What to do till the computer scientist comes. American Mathematical Monthly, 75(5):454–462, May 1968. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Forsythe:1969:CSE

[For69a] G. E. Forsythe. Computer science and education. In Morrell [Mor69], pages 1025–1039. ISBN 0-7204-2032-6. LCCN QA76 .I578.

Forsythe:1969:DTN

[For69b] George E. Forsythe. Design—then and now. In The Digest Record of the ACM-SIAM-IEEE 1969 Joint Conf. on Mathematical and Computer Aids to Design, pages 2–10. ACM Press, New York, NY 10036, USA, 1969. LCCN ????

Forsythe:1969:LDA

[For69c] George E. Forsythe. Let’s not discriminate against good work in design or experimentation. In AFIPS SJCC ’69 [AFI69], pages 538– 539. LCCN TK7885.A1 J6 1969.

Forsythe:1969:RPD

[For69d] George E. Forsythe. Remarks on the paper by Dekker. In Dejon and Henrici [DH69], pages 49–51. ISBN 0-471-20300-9. LCCN QA212 .C65.

Forsythe:1969:SQE

[For69e] George E. Forsythe. Solving a quadratic equation on a computer. In Boehm [Boe69], pages 138–152. LCCN QA11 .M39. Edited by the National Research Council’s Committee on Support of Research in the Mathematical Sciences (COSRIMS) with the collaboration of George A. W. Boehm.

Forsythe:1969:WSQ

[For69f] George E. Forsythe. What is a satisfactory quadratic equation solver? In Dejon and Henrici [DH69], pages 53–61. ISBN 0-471-20300-9. LCCN QA212 .C65.

Forsythe:1970:SQE

[For70a] G. Forsythe. Syllabi and qualifying examinations for the Ph.D. in computer science at Stanford University (continued). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(1):19–28, March 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Forsythe:1970:SEQ

[For70b] G. E. Forsythe. Stanford exam question. ACM SIGNUM Newsletter, 5(2):16, August 1970. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Forsythe:1970:MMP

[For70c] George E. Forsythe. The maximum and minimum of a positive definite quadratic polynomial on a sphere are convex functions of the radius. SIAM Journal on Applied Mathematics, 19(3):551–554, November 1970. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Forsythe:1970:PCWa

[For70d] George E. Forsythe. Pitfalls in computation, or why a math book isn’t enough. Technical Report CS 147, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1970. 43

pp. URL http://i.stanford.edu/pub/cstr/reports/cs/tr/70/
147/CS-TR-70-147.pdf.

Forsythe:1970:PCWb

[For70e] George E. Forsythe. Pitfalls in computation, or why a math book isn’t enough. American Mathematical Monthly, 77(9):931–956, November 1970. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Forsythe:1970:PEP

[For70f] George E. Forsythe. Problems, exams, projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(5): 37–43, December 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Forsythe:1971:F

[For71a] George E. Forsythe. Foreword. In Numerical Initial Value Problems in Ordinary Differential Equations [Gea71], page vii. ISBN 0-13-626606-1 (hardcover). LCCN QA372 .G4.

Forsythe:1971:RRS

[For71b] George E. Forsythe. Recent references on solving elliptic partial differential equations by finite differences or finite elements. ACM SIGNUM Newsletter, 6(1):99, 32–56, January 1971. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Forsythe:1972:NCMa

[For72a] George E. Forsythe. Von Neumann’s comparison method for random sampling from the normal and other distributions. Technical Report CS-TR-72-254, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1972. 21 pp.

Forsythe:1972:NCMb

[For72b] George E. Forsythe. Von Neumann’s comparison method for random sampling from the normal and other distributions. Mathematics of Computation, 26(120):817–826, October 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/journals/mcom/1972-26-120/S0025-
5718-1972-0315863-9/S0025-5718-1972-0315863-9.pdf.

Forsythe:19xx:NME

[Forxxa] George E. Forsythe. Numerical methods for elliptic partial differential equations. Report ????, Institute for Numerical Analysis, National Bureau of Standards, Los Angeles, CA, USA, ???? 19xx.

Forsythe:19xx:TSM

[Forxxb] George E. Forsythe. Theory of selected methods of finite matrix inversion and decomposition. Report ????, INA/UCLA, Los Angeles, CA, USA, ???? 19xx.

Forsythe:1958:CSN

[FR58a] George E. Forsythe and Paul C. Rosenbloom. Contemporary state of numerical analysis. Surveys in Applied Mathematics, 5:3–42, 1958. Issue on Numerical Analysis and Partial Differential Equations.

Forsythe:1958:NAP

[FR58b] George E. Forsythe and Paul C. Rosenbloom. Numerical analysis and partial differential equations, volume 5 of Surveys in Applied Mathematics. John Wiley and Sons, Inc. and Chapman and Hall, Ltd., New York, NY, USA and London, UK, 1958. x + 204 pp.

Forsythe:1942:RRM

[FS42] G. E. Forsythe and A. C. Schaeffer. Remarks on regularity of methods of summation. Bulletin of the American Mathematical Society, 48(??): 863–865, 1942. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-

881X (electronic). URL http://projecteuclid.org/euclid.bams/
1183504863.

Forsythe:1955:BCM

[FS55a] George E. Forsythe and E. G. Straus. On best conditioned matrices. Proceedings of the American Mathematical Society, 6(3):340– 345, June 1955. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-

6826 (electronic). URL https://www.ams.org/journals/proc/
1955-006-03/S0002-9939-1955-0069585-4.

Forsythe:1955:SFC

[FS55b] George E. Forsythe and Louise W. Straus. The Souriau–Frame characteristic equation algorithm on a digital computer. Journal of mathematics and physics / Massachusetts Institute of Technology, 34(??): 152–156, 1955. CODEN JMPHA9. ISSN 0097-1421.

Forsythe:1960:SPE

[FS60] George E. Forsythe and G. Szegő. Solution to problem E1398. American Mathematical Monthly, 67(??):696–697, ???? 1960. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Forsythe:1962:VDA

[FvT62] G. E. Forsythe, J. von der Groeben, and J. G. Toole. Vectorcardiographic diagnosis with the aid of ALGOL. Communications of the Association for Computing Machinery, 5(2):118–122, February 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1960:FDM

[FW60] George E. Forsythe and Wolfgang R. Wasow. Finite-Difference Methods for Partial Differential Equations. Applied Mathematics Series. John Wiley and Sons, Inc., New York, NY, USA, 1960. x + 444 pp.

Forsythe:1965:AGPa

[FW65a] George E. Forsythe and Niklaus Wirth. Automatic grading programs. Technical Report CS-TR-65-17, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1965. ii + 17

pp. URL http://i.stanford.edu/pub/cstr/reports/cs/tr/65/
17/CS-TR-65-17.pdf.

Forsythe:1965:AGPb

[FW65b] George E. Forsythe and Niklaus Wirth. Automatic grading programs. Communications of the Association for Computing Machinery, 8(5): 275–278, May 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gear:1971:NIV

[Gea71] Charles William Gear. Numerical Initial Value Problems in Ordinary Differential Equations. Prentice-Hall series in automatic computation. Prentice-Hall, Inc., Englewood Cliffs, NJ 07632, USA, 1971. ISBN 0-13-626606-1 (hardcover). xvii + 253 pp. LCCN QA372 .G4.

Herriot:1972:MGF

[Her72] John G. Herriot. In memory of George E. Forsythe. Communications of the Association for Computing Machinery, 15(8):719–720, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles in honor of George E. Forsythe (ACM Student Competition Award Papers).

Holmboe:1945:DMC

[HFG45] Jörgen Holmboe, George E. (George Elmer) Forsythe, and William Sharp Gustin. Dynamic meteorology. John Wiley and Sons, Inc., New York, NY, USA, 1945. xvi + 378 pp. LCCN QC861 .H6.

Householder:1951:MCM

[HFG51] Alston S. Householder, George E. Forsythe, and Hallett-Hunt Germond, editors. Monte Carlo method. Proceedings of a Symposium Held June 29, 30 and July 1, 1949 in Los Angeles, California, volume 12 of Applied Mathematics Series / National Bureau of Standards. United States Government Printing Office, Washington, DC, USA, 1951.

Householder:1973:GFJ

[Hou73] A. S. Householder. George E. Forsythe (January 8, 1917–April 9, 1972). SIAM Journal on Numerical Analysis, 10:viii–xi, 1973. CO-DEN SJNAAM. ISSN 1095-7170. Collection of articles dedicated to the memory of George E. Forsythe.

Knuth:1972:GFD

[Knu72] Donald E. Knuth. George Forsythe and the development of Computer Science. Communications of the Association for Computing Machinery, 15(8):721–726, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles in honor of George E. Forsythe (ACM Student Competition Award Papers).

Knuth:1996:GFD

[Knu96a] Donald E. Knuth. George Forsythe and the development of Computer Science. In Selected Papers on Computer Science [Knu96b], chapter 14, pages 241–259. ISBN 1-881526-91-7 (paperback), 1-881526-92-5 (hardcover). LCCN QA76.K537 1996. Reprint of [Knu72]. Includes Forsythe publication list.

Knuth:1996:SPC

[Knu96b] Donald Ervin Knuth. Selected Papers on Computer Science, volume 59 of CSLI lecture notes. Cambridge University Press, Cambridge, UK, 1996. ISBN 1-881526-91-7 (paperback), 1-881526-92-5 (hardcover). xii + 274 pp. LCCN QA76.K537 1996. Dedicated “to George Forsythe, whose inspiring leadership shaped the field.”.

Lanczos:1952:SSL

[Lan52] Cornelius Lanczos. Solution of systems of linear equations by minimized-iterations. Journal of Research of the National Bureau of Standards (1934), 49:33–53, 1952. ISSN 0091-0635.

Lee:1973:VSN

[LF73] E. H. Lee and G. E. Forsythe. Variational study of nonlinear spline curves. SIAM Review, 15(1):120–133, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Misa:2017:CCC

[Mis17] Thomas J. Misa, editor. Communities of computing: computer science and society in the ACM, volume 13 of ACM books. ACM Press, New York, NY 10036, USA, 2017. ISBN 1-970001-84-4. ISSN 2374-6777. xiii + 408 pp. LCCN QA76.25 .C6498 2017.

Moler:1972:MGF

[Mol72] Cleve Moler. A memory of George Forsythe. ACM SIGNUM Newsletter, 7(3):8–9, October 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Morrell:1969:IPP

[Mor69] A. J. H. Morrell, editor. Information processing 68: proceedings of IFIP congress 1968, organized by the International Federation for Information Processing, Edinburgh, 5–10 August 1968. North-Holland, Amsterdam, The Netherlands, 1969. ISBN 0-7204-2032-6. LCCN QA76 .I578.

Nash:1990:HSC

[Nas90] Stephen G. Nash, editor. A History of Scientific Computing. ACM Press history series. Addison-Wesley and ACM Press, Addison-Wes-ley and New York, NY 10036, USA, 1990. ISBN 0-201-50814-1. xix + 359 pp. LCCN QA76.17 .H59 1990.

November:2017:GFC

[Nov17] Joseph November. George Forsythe and the creation of computer science as we know it. In Misa [Mis17], chapter 3, pages 47–70. ISBN 1-970001-84-4. ISSN 2374-6777. LCCN QA76.25 .C6498 2017.

Parlett:1973:GTP

[PP73] B. N. Parlett and W. G. Poole, Jr. A geometric theory for the QR, LU and power iterations. SIAM Journal on Numerical Analysis, 10(2):389–412, April 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https://www.math.utah.edu/

pub/tex/bib/gvl.bib;   https://www.math.utah.edu/pub/tex/

bib/siamjnumeranal.bib. Collection of articles dedicated to the memory of George E. Forsythe.

Paige:1953:SLE

[PT53] L. J. (Lowell J.) Paige and Olga Taussky, editors. Simultaneous Linear Equations and the Determination of Eigenvalues[: proceedings of a symposium held 23–25 August 1951 at the Institute for Numerical Analysis, Los Angeles, California], volume 29 of Applied Mathematics Series. U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1953. LCCN QA3 .U5 no. 29.

Rosenbloom:1960:RPN

[RFW60] Paul C. Rosenbloom, George E. Forsythe, and Lawrence A. Weller. Recent publications: Numerical analysis and partial differential equations. volume V. surveys in applied mathematics. American Mathematical Monthly, 67(3):306, 1960. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Read:1972:NGR

[RH72] Ronald C. Read and K. Harada. A note on the generation of rosary permutations. with a response by K. Harada. Communications of the Association for Computing Machinery, 15(8):775, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles in honor of George E. Forsythe (ACM Student Competition Award Papers).

Spjotvoll:1972:NTF

[Spj72] Emil Spjotvoll. A note on a theorem of Forsythe and Golub. SIAM Journal on Applied Mathematics, 23(3):307–311, November 1972. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Taussky:1954:CSS

[Tau54] Olga Taussky, editor. Contributions to the Solution of Systems of Linear Equations and the Determination of Eigenvalues, volume 39 of Applied Mathematics Series. U.S. National Bureau of Standards, Gaithersburg, MD, USA, September 30, 1954.

Taub:1961:JNCa

[Tau61a] A. H. Taub, editor. John von Neumann: Collected Works: Volume I: Logic, Theory of Sets and Quantum Mechanics. Pergamon Press, New York, NY, USA, 1961. x + 654 pp. LCCN ???? See also volumes II–VI [Tau61b, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1961:JNCb

[Tau61b] A. H. Taub, editor. John von Neumann: Collected Works. Volume II: Operators, Ergodic Theory and Almost Periodic Functions in a Group. Pergamon Press, New York, NY, USA, 1961. x + 568 pp. LCCN ???? See also volumes I, III–VI [Tau61a, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1962:JNC

[Tau62] A. H. Taub, editor. John von Neumann: Collected Works. Volume IV: Continuous Geometry and Other Topics. Pergamon Press, New York, NY, USA, 1962. x + 516 pp. LCCN ???? See also volumes I–III, V–VI [Tau61a, Tau61b, Tau63a, Tau63b, Tau63c].

Taub:1961:JNCc

[Tau63a] A. H. Taub, editor. John von Neumann: Collected Works. Volume III: Rings of Operators. Pergamon Press, New York, NY, USA, 1961– 1963. ix + 574 pp. LCCN ???? See also volumes I–II, IV–VI [Tau61a, Tau61b, Tau62, Tau63b, Tau63c].

Taub:1963:JNCa

[Tau63b] A. H. Taub, editor. John von Neumann: Collected Works. Volume V: Design of Computers, Theory of Automata and Numerical Analysis. Pergamon Press, New York, NY, USA, 1963. ix + 784 pp. LCCN ???? See also volumes I–IV, VI [Tau61a, Tau61b, Tau63a, Tau62, Tau63c].

Taub:1963:JNCb

[Tau63c] A. H. Taub, editor. John von Neumann: Collected Works. Volume VI: Theory of Games, Astrophysics, Hydrodynamics and Meteorology. Pergamon Press, New York, NY, USA, 1963. x + 538 pp. LCCN ???? See also volumes I–V [Tau61a, Tau61b, Tau63a, Tau62, Tau63b].

Varah:1987:WGF

[Var87] J. Varah. The work of George Forsythe and his students. In Crane [Cra87], pages 139–150. ISBN 0-89791-229-2. LCCN QA76 .A25 1987.

Varah:1990:IGF

[Var90] James M. Varah. The influence of George Forsythe and his students. In Nash [Nas90], pages 31–40. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Vazov:1963:RMR

[VF63] V. Vazov and Dž. Forsaĭt. Raznostnye metody resheniya differentsialnykh uravnenii v chastnykh proizvodnykh. (Russian) [Finite-Difference Methods for Partial Differential Equations]. Izdatel’stvo Inostran. Lit., Moscow, USSR, 1963. 487 pp. Translated from the English by B. M. Budak and N. P. Židkov.

vonNeumann:1951:VTU

[vN51] John von Neumann. 13. Various techniques used in connection with random digits. In Householder et al. [HFG51], pages

36–38. URL https://dornsifecms.usc.edu/assets/sites/520/
docs/VonNeumann-ams12p36-38.pdf. Summary written by G. E.

Forsythe. Reprinted in [Tau63b, Paper 23, pp. 768–770].

Weiss:1988:BOP

[Wei88] Eric A. Weiss. Biographies: Oh, pioneers! Annals of the History of Computing, 10(4):348–361, October/December 1988. CO-DEN AHCOE5. ISSN 0164-1239. URL http://dlib.computer.

org/an/books/an1988/pdf/a4348.pdf;   http://www.computer.
org/annals/an1988/a4348abs.htm.